why 0! is one. and what the actual defination of factorial how it can be justified that 1! is also one.

Grade:12th pass

1 Answers

Vijay Mukati
7 years ago
Since, n! = n (n-1)!, so from here for n = 1, 1! = 1 * 0! and thus 0! = 1.

Thanks, Vj